(a) worked out incorrectly; or (b) worked out correctly but using a method not specified in the problem statement; or (c) not5) A partial sum of an arithmetic sequence is given, find the sum. (5 points) 4 + 7 + 10 + ... + 301
Answer #1

4 + 7 + 10+ 301 a +9₂ +93 + an On = a, + (n-1)d d=9₂-9,=7-4=3 (A =640 an=301= 4+ (n-1)3 301-4=(-1)3 da 297 = (n-1)3 OPS 20d n